OpenStudy (anonymous):

If ZW/RT = 6/5 and the area of triangle QRS = 75cm^2, how do I find the area of triangle XYZ?

OpenStudy (anonymous):

Here is the figure :p

OpenStudy (anonymous):

#1 eww

OpenStudy (anonymous):

Eww?

OpenStudy (anonymous):

Area of triangle = 1/2 b.h for QRS area is 75cm sq and height is 5 75=1/2*5*b b=30cm since the triangles have the same angles the proportions are the same. XYZ dimensions= base= 30*(6/5) = 36 cm height already known at 6cm area = 1/2 (6)(36)= 108

OpenStudy (anonymous):

the question i mean lol

OpenStudy (anonymous):

Oh yeah >.< I hate geometry.. Thx for the help tho love :)

OpenStudy (anonymous):

lol ya i dont care for geometry or math tbh lmao and np ;P